An electric power plant uses solid waste for fuel in the production of electricity. The cost Y in dollars per hour to produce electricity is Y = 12 + 0.3X + 0.27X2, where X is in megawatts. Revenue in dollars per hour from the sale of electricity is 15X−0.2X2. Find the value of X that gives maximum profit


1
Expert's answer
2021-10-03T14:17:03-0400

"Cost = Y = 12 +0.3X+0.27X^2 \\
Revenue = R = 15X -0.2X^2 \\
Profit = R-Y \\
Π = 15X- 0.2X^2 -12 -0.3X -0.27X^2 \\
\frac{dΠ}{dX} = 15 -0.4X -0.3 -0.54X = 0 \\
14.7 -0.94X = 0 \\
X = \frac{14.7}{0.94} = 15.64 \;megawatts"